OpenStudy (anonymous):

consider the following function and simplify the expression. f(x+h)-f(x) divided by h. f(x)=2x+4 Please help me. I need this for more my Calculus class.

OpenStudy (anonymous):

it is just the slope what is the slope of \(y=2x+4\) ?

OpenStudy (anonymous):

It's 2 right? But what about the function?

OpenStudy (anonymous):

it is a linear function the change in y over the change in x is a constant, namely the slope

OpenStudy (anonymous):

go ahead and compute it \[f(x+h)=2(x+h)+4=2x+2h+4\] \[f(x+h)-f(x)=2x+2h+4-(2x+4)=2h\] divide by \(h\) and get the slope

OpenStudy (anonymous):

That was right. I enter my homework online, so I know right away whether it is right or wrong. I still don't understand how you got the answer. thanks a lot though

OpenStudy (anonymous):

i didn't skip any steps see if you can understand what i wrote